OpenStudy (anonymous):

find cartesian equation of the curve traced out by x=a cos^3(theta), y=a sin^3(theta)

hartnn (hartnn):

ok, do u have any idea how to start ? have u tried ?

hartnn (hartnn):

i would like to give you hint: use \(\large \sin^2 t+\cos^2t=1\)

OpenStudy (anonymous):

I got to cos^2(theta)+sin^2(theta)=x/(acostheta)+y/(asintheta)

OpenStudy (anonymous):

thus 1=x/asec(theta)+y/a csc(theta)

hartnn (hartnn):

basically, u need to eliminate theta from those 2 equations, right ?

OpenStudy (anonymous):

yes but I'm at a loss now

hartnn (hartnn):

thats not cartesiian form...

OpenStudy (anonymous):

Once you get rid of theta then it's cartesian right? I just don't know how to get rid of sec and csc now

hartnn (hartnn):

right. ok, let me give you another hint : x/a = cos^3 theta so, (x/a)^(2/3) = cos^2 theta did u get this ?

OpenStudy (anonymous):

ok... but if i substitute that into 1=(x/a)sectheta+(y/a)csctheta, I get 1=cos^3 theta/ cos theta + sin^3 theta/sin theta, so basically 1=cos^2 theta + sin^2 theta, which is 1=1...

hartnn (hartnn):

don't do that! i found cos^2 theta =.... can u find sin^2 theta =... ?

OpenStudy (anonymous):

do you mean cos^2 theta= x/(a cos theta)

hartnn (hartnn):

no! x/a = cos^3 theta so, (x/a)^(2/3) = cos^2 theta did u get this ?

hartnn (hartnn):

i just raised the power of 2/3 on both sides to convert cos^3 theta to cos^2 theta....

OpenStudy (anonymous):

ok but where am I supposed to substitute that in for?

hartnn (hartnn):

u got cos^2 theta as (x/a)^(2/3) right ? now can u find, similarly, sin^2 theta ? in terms of y and a ?

OpenStudy (anonymous):

it's (y/a)^(2/3)

hartnn (hartnn):

correct. now just add the 2 equations and you would have successfully eliminated the theta

hartnn (hartnn):

just because cos^2 theta +sin^2 theta =1

OpenStudy (anonymous):

ohhh ok

OpenStudy (anonymous):

would the simplest answer then be 1=(x/a)^(2/3)+(y/a)^(2/3)

hartnn (hartnn):

yes, exactly.

OpenStudy (anonymous):

ok got it THANK YOU SO MUCH!!!!

hartnn (hartnn):

\[ \begin{array}l\color{red}{\text{w}}\color{orange}{\text{e}}\color{#e6e600}{\text{l}}\color{green}{\text{c}}\color{blue}{\text{o}}\color{purple}{\text{m}}\color{purple}{\text{e}}\color{red}{\text{ }}\color{orange}{\text{^}}\color{#e6e600}{\text{_}}\color{green}{\text{^}}\color{blue}{\text{}}\end{array} \]